# Learning relevant contextual variables within Bayesian optimization

### v3

@@ 3.2 Sensitivity-Analysis-Driven CBO (SADCBO)
Building on top of the variable selection method of Section 3.1, we now present SADCBO, a sequential method for performing BO in the presence of irrelevant contextual variables, summarized in Algorithm 1. SADCBO proceeds in two phases.
@@
The information gained from purely observing contextual variables will, however, saturate at some point, leading to diminishing simple regret differences. This motivates the introduction of a second phase, in which contextual variables can have their values arbitrarily set, through optimization.
In the second, contextual optimization phase, we optimize the contextual variables selected at each iteration based on their FC relevance. As optimizing a context variable $z^{(j)}$ is associated with a cost $\lambda_{j}$, we modify the FC relevance in Equation 5 to be $\text{FC}(j)/\lambda_{j}$. Our variable selection criterion can then be interpreted as the degree of sensitivity per unit cost. This allows SADCBO to automatically trade off a variable’s potential to substantially affect the optimum with the associated optimization cost. As before, once the contextual variables $\mathbf{z}^{(J_{\eta})}$ have been selected, we train a GP surrogate based on $\{(\mathbf{x}_{i},\mathbf{z}_{i}^{(J_{\eta})},\mathbf{y}_{i})\}_{i=1}^{t}$ and select the next design and contextual variables to query as
@@
In effect, $J_{\eta}$ represents our approximation for $J^{\star}$ as introduced in Equation 3. Note that our acquisition function is not cost-weighted, as cost-weighted acquisition functions can dramatically underperform (Lee et al., 2021), specifically for non-continuous cost models. Including the cost at the model selection level avoids this issue.
Switching from observational to contextual optimization phase.
We use the criterion proposed by Ishibashi et al. (2023) for determining the stopping time in BO. Using this criterion, we detect the point at which the optimization gain based on purely observing the contextual variables diminishes, following which the contextual optimization phase begins. We now briefly describe the details of this switching criterion.
With $\mathbf{v}=[\mathbf{x},\mathbf{z}]$, let $\mathbf{v}^{\star}_{t}=\operatorname*{arg\,max}_{\mathbf{v}\in\mathcal{D}_{t}}f(\mathbf{v})$ be the current best candidate point in the dataset up to time $t$. Denoting $f^{\star}:=\max_{\mathbf{v}\in\mathcal{V}}f(\mathbf{v})$, let $R_{t}=f^{\star}-\mathbb{E}_{\hat{f}\sim p(f|\mathcal{D}_{t})}[\max_{\mathbf{v}\in\mathcal{V}}~{}\hat{f}(\mathbf{v})]$ be the expected minimum simple regret. Then, with probability $1-\delta$, $\Delta R_{t}=\lvert R_{t}-R_{t-1}\rvert$ can be upper bounded by $\Delta\tilde{R}_{t}$ with
@@
Further details about the derivation of $s_{t}$ and the expression of $\kappa_{\delta,t-1}$ can be found in Appendix A.
@@
6:     Compute $\text{FC}_{\mathcal{D}_{t}^{\text{BO}}}(j)$ for all $j$ (Equation 4)
@@
16:     $\Lambda\leftarrow\Lambda-\lambda_{\mathbf{x}}+\sum_{j\in J_{\eta}}\lambda_{j},~{}t\leftarrow t+1$
@@ 4 Related Work
Bogunovic et al. (2018); Kirschner et al. (2020); Husain et al. (2023); Saday et al. (2023) perform worst-case optimization under fluctuations of the contextual variables. In particular, Distributionally-Robust BO (DRBO, Kirschner et al., 2020) tries to maximize the expected black-box function value under the worst-case distribution of the contextual variables. This worst-case distribution belongs to an “uncertainty set”, a ball centered around a reference distribution that is gradually learned (Tulabandhula and Rudin, 2014). However, as in Krause and Ong (2011), these works assume that the relevant contextual variables are known a priori, and can only be observed, after the designs have been selected, and not controlled.
@@
Due to the curse of dimensionality, the performance of standard BO is severely degraded when applied in high-dimensional input spaces. To tackle this problem, most proposed approaches either aim at carrying out BO in a lower-dimensional space instead of the original or work with a structured GP surrogate. A lower-dimensional subspace can be found in a data-agnostic manner, for instance by randomly dropping dimensions of the problem (Li et al., 2017) or considering tree-like random decompositions (Ziomek and Bou-Ammar, 2023). Data-driven methods based on various measures of feature relevance have also been proposed (Spagnol et al., 2019; Shen and Kingsford, 2021). In contrast, structured surrogate methods encode structural information about the objective, for instance using an additive kernel, together with an acquisition function which is additive under the provided decomposition (Rolland et al., 2018). Finally, Eriksson and Jankowiak (2021) and Liu et al. (2023) proposed using a sparsity-enforcing GP surrogate, equipped with a heavy-tailed horseshoe prior on the squared inverse lengthscales.
@@
In most methods, the BO budget is given in iterations, implicitly assuming that each evaluation has the same cost. In practice, cost may vary significantly across different regions of the input space (Lee et al., 2020), or depend on the number of variables we optimize over. Cost-aware BO integrates the cost-constrained nature of the problem, usually within the acquisition function, but also with more involved strategies like constrained Markov decision processes when the total budget is known beforehand (Lee et al., 2021). The recent work by Tay et al. (2023) carries out Robust BO while at the same time involving a notion of controlled variables at a cost. However—unlike our framework—they require the nonselected variables to be sampled from a known distribution at each iteration.
@@ 5 Experimental Results
We evaluate our approach on several real-world examples and synthetic functions, described in Section 5.1. We compare against multiple baselines (Table 1) and present results in Section 5.2. In Section 5.3 we discuss the influence of various experimental settings: number of noise variables present, contextual variable query cost, surrogate and method hyperparameters. Code will be available upon acceptance.
Baselines.
We benchmark our approach, coined SADCBO, against baselines referenced in Table 1. In particular, MMDBO operates variable selection in BO through an MMD-based measure of sensitivity (Spagnol et al., 2019) and is detailed in Appendix B, whereas Dropout (Li et al., 2017) randomly selects half of the contextual variable for optimization. Next, CaBO (Lee et al., 2020) performs vanilla BO over $[\mathbf{x},\mathbf{z}]$, using a cost-weighted acquisition function. The cost model employed here is a smoothed version of our noncontinuous cost model, using a Gaussian curve. Finally, CBO refers to the Contextual BO framework proposed by Krause and Ong (2011). As a way to assess the impact of contextual variables and selection mechanisms, we also report CUBO and VBO: Context-Unaware BO over the designs $\mathbf{x}$ only, and Vanilla BO over both design and contextual variables $[\mathbf{x},\mathbf{z}]$.
@@
CaBO Cost-Aware BO over $[\mathbf{x},\mathbf{z}]$ (Lee et al., 2020)
CBO Contextual BO using all contexts $\mathbf{z}$ (Krause and Ong, 2011)
With variable selection Dropout Randomly drop half of context variables (Li et al., 2017)
MMDBO Maximum mean discrepancy-driven BO (Spagnol et al., 2019)
@@
Implementation details.
We fix the hyperparameter of SADCBO to $\eta=0.8,Q=10,\gamma_{t}=0.8~{}\forall t$. For the GP surrogate, we use a squared-exponential kernel with independent lengthscales for each variable, learned through marginal likelihood maximization. We use the UCB acquisition strategy, as well as $Q$-UCB for computing $\mathcal{D}_{t}^{Q}$ (Equation 7(Wilson et al., 2017). In all experiments, we assume that all variables, design or contextual ones, have cost $\lambda_{j}=1~{}\forall j\in\{1,\dots,d+c\}$, except in a dedicated experiment in Section 5.3. Our algorithm is implemented using the BoTorch framework (Balandat et al., 2020).
@@ 5.1 Experiments
We benchmark on 4 real-world and 4 synthetic experiments (Table 2) described in brief here and detailed in Appendix D.
Table 2: Dimensionality of the experiments. For synthetic experiments, n.v. stands for (artificial) noise variables, added on top of the design and contextual variables.
@@
EggHolder 2 + 4 n.v. 1 1
Hartmann4D 4 + 3 n.v. 2 2
Hartmann6D 6 + 6 n.v. 3 3
Ackley 5 + 8 n.v. 2 3
Portfolio Optimization 5D.
This dataset was first introduced by Cakmak et al. (2020). The goal is to optimize three design variables, which stand for the hyperparameters of a trading strategy, to maximize return under random environmental conditions. There are two contextual variables, namely: bid–ask spread and the borrowing cost.
Yacht Hydrodynamics 6D.
This dataset comes from the UCI Machine Learning Repository (Gerritsma et al., 2013). The optimization problem is to maximize the residuary resistance per unit weight of displacement of a yacht by controlling its 5-dimensional hull geometry coefficients. Design variables the first four dimensions of the hull geometry coefficients. The contextual variables are the last hull geometry dimension and the Froude number.
Molecule structure optimization 21D.
This computational chemistry example consists of optimizing the bond angles in an alanine molecule to determine the lowest energy conformer, i.e., the structure the molecule will likely take in nature. These problems are complicated by high dimensionality. Alanine has 21 angular variables: 3 key variables based on prior domain knowledge set as design variables, and 18 other angles treated as contextual variables. Molecular energies are calculated with the AMBER forcefield (Case et al., 2023) at each round of BO.
Robot pushing task 14D.
We follow Wang et al. (2017) and consider a control parameter tuning problem for robot pushing. This real-world function returns the distance between a designated goal location and two objects being pushed by two robot hands, whose trajectory is determined by 14 parameters specifying the location, rotation, velocity and moving direction, among others. There are 6 design variables and 8 contextual variables.
Synthetic experiments.
We also consider four synthetic test functions, see Tables 2 and D.2 for details. A min-max transformation is performed on the input data, scaling it to the unit cube: $\mathcal{X}\times\mathcal{Z}=[0,1]^{d+c}$. Similarly, the output is scaled between $[0,1]$ and a noise term $\varepsilon\sim\mathcal{N}(0,\sigma^{2}_{\text{noise}})$ is added with $\sigma^{2}_{\text{noise}}=0.001$. The contextual variable distribution is $p(\mathbf{z})=\mathcal{U}([0,1]^{c})$.
Figure 2: Benchmark of the different methods. (a) On real-world datasets, SADCBO (red curve with white markers) performs on par with other baselines and is the top performer for the Robot Pushing task. (b) On synthetic functions, SADCBO outperforms other baselines in three cases out of four. (c) Histograms of early stopping criterion hitting time for SADCBO computed for the Hartmann6D (c.1) and Hartmann4D problems (c.2). (d) Inclusion probability of each contextual variable for SADCBO computed for the Hartmann6D (d.1) and Hartmann4D problems (d.2). Each panel shows the mean ±$\pm 2$ standard error across $N=100$ trials.
@@ 5.2 Results
In each plot from Figure 2, we report the best value found by each baseline as a function of the number of iterations. In real-world experiments (Figure 2a), SADCBO (in red with white markers) quickly converges to the optimum. SADCBO consistently outperforms the first baselines VBO and CUBO, even though in the Molecular Shape example, SADCBO and CUBO perform on par due to the good choices of the domain experts on the design variables. Except for the Robot Pushing task, the difference between SADCBO and CBO (in blue) is marginal in the real-world experiments. The latter enhances the surrogate model with information from sampled contexts, while our method may even optimize selected contextual variables if needed. Given that these baselines perform similarly, combined with the observation that optimizing only design variables (CUBO, in yellow) produces poor results for the Portfolio and Yacht problems, we can conclude that contextual variables play a significant part in maximizing these two objectives. The cost-aware BO baseline CaBO performs poorly in all tasks. Dropout and MMDBO consistently underperforms, except on the Yacht example for the latter. These baselines operate variable selection in a random manner for Dropout and using Hilbert-Schmidt Independence Criterion for MMDBO (Gretton et al., 2007), two strategies that do not seem to surpass the Feature Collapse method implemented in SADCBO. This observation highlights the need for an informed variable selection strategy. In-depth findings for the Molecule experiment are presented in Section D.1 and provide additional explanations as to why SADCBO clearly outperforms MMDBO and Dropout.
@@
Figure 2b displays the best value found by each baseline for synthetic functions. SADCBO ranks first on 3 out of 4 examples, closely followed by the cost-aware baseline CaBO, which performs much better on synthetic experiments than on the real-world ones. The contextual BO baseline CBO that obtained second to best results in real-world experiments, is now less performant, due to the fact that it does not optimize the context, like CUBO. This seems to be particularly critical for Ackley5D, whereas for Hartmann6D/Hartmann4D, simply enhancing the surrogate with contextual variable observation already leads to a large performance gap between CUBO and CBO. Lastly, VBO does a poor job as it optimizes every variable, thus spending a large fraction of the budget every iteration.
For Hartmann6D and Hartmann4D, Figure 2c reports the time at which SADCBO’s stopping criterion (Equation 11) kicks in, in proportion to the total budget, demonstrating that both phases are leveraged in our approach. We ensure that the criterion is well-behaved: the more information about the output is contained in the contextual variables, the later the early stopping criterion triggers (Figure S1). Even though the stopping criterion was initially devised for vanilla BO, its application in a CBO setting is fruitful.
Lastly, Figure 2d reports the sensitivity indices computed at each iteration for each contextual variable, averaged across whole trajectories of multiple trials. For Hartmann6D, the results match the Sobol sensitivity analysis results (Table S1), even though global sensitivity indices may differ from sensitivity indices with respect to the function optimum. Similar findings apply to Hartmann4D (Table S2). Results for other problems can be found in Figure S2.

### v4

@@ 3.2 Sensitivity-Analysis-Driven CBO (SADCBO)
Building on top of the variable selection method discussed in Section 3.1, we now present SADCBO, a sequential method for performing BO in the presence of irrelevant contextual variables. SADCBO proceeds in two phases.
@@
In the early stage of the optimization, cheap queries where contextual variables are not optimized still provide a considerable amount of information. The information gained from purely observing contextual variables will, however, saturate at some point, leading to diminishing simple regret differences. At this point, it becomes necessary to pay the higher price to control more dimensions of the input space. This motivates the introduction of a second phase, in which contextual variables can have their values arbitrarily set, through optimization.
In the second, contextual optimization phase, we optimize the contextual variables selected at each iteration based on their FC relevance. As optimizing a context variable $z^{(j)}$ is associated with a cost $\lambda_{j}$, we modify the FC relevance in Equation 5:
$\tilde{\text{FC}}_{\mathcal{D}_{t}}(j)=\text{FC}_{\mathcal{D}_{t}}(j)/\lambda_{j}$ (9)
Our variable selection criterion can then be interpreted as the degree of sensitivity per unit cost. This allows SADCBO to automatically trade off a variable’s potential to greatly affect the optimum with the associated optimization cost. As before, once the contextual variables $\mathbf{z}^{(J_{\eta})}$ have been selected, we train a GP surrogate using $\{(\mathbf{x}_{i},\mathbf{z}_{i}^{(J_{\eta})},y_{i})\}_{i=1}^{t}$ and select the next design and contextual variables to query as
@@
In effect, $J_{\eta}$ represents our approximation for $J^{\star}$ as introduced in Equation 3. Note that our acquisition function is not cost-weighted, as cost-weighted acquisition functions can dramatically underperform [Lee et al., 2021], specifically for non-continuous cost models. Including the cost at the model selection level avoids this issue.
Switching from observational to optimization phase.
We employ the criterion proposed by Ishibashi et al. [2023] for determining the stopping time in BO. Using this criterion, we detect the point at which the optimization gain based on purely observing the contextual variables diminishes, following which the contextual optimization phase begins. We now briefly describe the details of this switching criterion.
With $\mathbf{v}=[\mathbf{x},\mathbf{z}]$, let $\mathbf{v}^{\star}_{t}=\operatorname*{arg\,max}_{\mathbf{v}\in\mathcal{D}_{t}}f(\mathbf{v})$ be the current best candidate point in the dataset up to time $t$. Denoting $f^{\star}:=\max_{\mathbf{v}\in\mathcal{V}}f(\mathbf{v})$, let $R_{t}=f^{\star}-\mathbb{E}_{\hat{f}\sim p(f|\mathcal{D}_{t})}[\max_{\mathbf{v}\in\mathcal{V}}\leavevmode\nobreak\ \hat{f}(\mathbf{v})]$ be the expected minimum simple regret. Then, with probability $1-\delta$, $\Delta R_{t}=\lvert R_{t}-R_{t-1}\rvert$ can be upper bounded by $\Delta\tilde{R}_{t}$ with
@@
Further details about the derivation of $s_{t}$ and the expression of $\kappa_{\delta,t-1}$ can be found in Appendix B. The entire algorithm is summarized in Algorithm 1.
@@
6:     Compute $\text{FC}_{\mathcal{D}_{t}^{\text{BO}}}(j)$ for all $j$ (Equation 5 or (9) if phase = optimization)
@@
16:     $\Lambda\leftarrow\Lambda-\lambda_{\mathbf{x}}+\sum_{j\in J_{\eta}}\lambda_{j},\leavevmode\nobreak\ t\leftarrow t+1$
@@ 4 Related work
Bogunovic et al. [2018], Kirschner et al. [2020], Husain et al. [2023] and Saday et al. [2023] perform worst-case optimization under fluctuations of the contextual variables. In particular, Distributionally-Robust BO [DRBO, Kirschner et al., 2020] tries to maximize the expected black-box function value under the worst-case distribution of the contextual variables. This worst-case distribution belongs to an “uncertainty set”, a ball centered around a reference distribution that is gradually learned [Tulabandhula and Rudin, 2014]. However, as in Krause and Ong [2011], these works assume that the relevant contextual variables are known a priori, and can only be observed, after the designs have been selected, and not controlled.
@@
Due to the curse of dimensionality, the performance of standard BO is severely degraded when applied in high-dimensional input spaces. To tackle this problem, most approaches either aim at carrying out BO in a lower-dimensional space instead of the original or work with a structured GP surrogate. A lower-dimensional subspace can be found in a data-agnostic manner, for instance by randomly dropping dimensions of the problem [Li et al., 2017] or considering tree-like random decompositions [Ziomek and Bou-Ammar, 2023]. Data-driven methods based on various measures of feature relevance have also been proposed [Spagnol et al., 2019, Shen and Kingsford, 2021]. In contrast, structured surrogate methods encode structural information about the objective, for instance using an additive kernel, yielding an acquisition function that is additive under the provided decomposition [Rolland et al., 2018]. Finally, Eriksson and Jankowiak [2021] and Liu et al. [2023] proposed using a sparsity-enforcing GP surrogate, equipped with a heavy-tailed horseshoe prior on the squared inverse lengthscales.
@@
In most methods, the BO budget is given in iterations, implicitly assuming that each evaluation has the same cost. In practice, cost may vary significantly across different regions of the input space [Lee et al., 2020], or depend on the number of variables we optimize over. Cost-aware BO integrates the cost-constrained nature of the problem, usually within the acquisition function. Let us also mention more involved strategies like constrained Markov decision processes when the total budget is known beforehand [Lee et al., 2021]. The recent work by Tay et al. [2023] carries out Robust BO while at the same time involving a notion of controlled variables at a cost. However—unlike our framework—they require the nonselected variables to be sampled from a known distribution at each iteration.
@@ 5 EXPERIMENTAL RESULTS
We evaluate our approach on several real-world examples and synthetic functions, described in Section 5.1. We compare against multiple baselines (Table 1) and present results in Section 5.2. In Section 5.3 we discuss the influence of various experimental settings: number of noise variables present, contextual variable query cost, surrogate and method hyperparameters. We conclude by presenting several insights regarding the phase-switching criterion.
Baselines. We benchmark our approach, coined SADCBO, against baselines referenced in Table 1. In particular, MMDBO operates variable selection in BO through an MMD-based measure of sensitivity [Spagnol et al., 2019] and is detailed in Appendix C, whereas Dropout [Li et al., 2017] randomly selects half of the contextual variable for optimization. Next, CaBO [Lee et al., 2020] performs vanilla BO over $[\mathbf{x},\mathbf{z}]$, using a cost-weighted acquisition function. The cost model employed here is a smoothed version of our noncontinuous cost model, using a Gaussian curve. Finally, CBO refers to the Contextual BO framework proposed by Krause and Ong [2011]. As a way to assess the impact of contextual variables and selection mechanisms, we also report CUBO and VBO: Context-Unaware BO over the designs $\mathbf{x}$ only and Vanilla BO over both design and contextual variables $[\mathbf{x},\mathbf{z}]$.
@@
CaBO Cost-Aware BO over $[\mathbf{x},\mathbf{z}]$ [Lee et al., 2020]
CBO Contextual BO using all contexts $\mathbf{z}$ [Krause and Ong, 2011]
With variable selection Dropout Randomly drop half of context variables [Li et al., 2017]
MMDBO Maximum mean discrepancy-driven BO [Spagnol et al., 2019]
@@
Implementation details. We fix the hyperparameters of SADCBO to $\eta=0.8,Q=10,\gamma_{t}=0.8\leavevmode\nobreak\ \forall t$. For the GP surrogate, we use a squared-exponential kernel with independent lengthscales for each variable, learned through marginal likelihood maximization. We use the UCB acquisition strategy, as well as $Q$-UCB for computing $\mathcal{D}_{t}^{Q}$ (Equation 7[Wilson et al., 2017]. In all experiments, we assume that any variable, design or contextual ones, has cost $\lambda_{j}=1\leavevmode\nobreak\ \forall j\in\{1,\dots,d+c\}$, except in a dedicated study in Section 5.3. Our algorithm is implemented using the BoTorch framework [Balandat et al., 2020]. Code can be accessed at https://github.com/julienmartinelli/SADCBO.
@@ 5.1 Experiments
We benchmark on 4 real-world and 4 synthetic experiments (Table 2) described in brief here and detailed in Appendix E.
Table 2: Dimensionality of the experiments. For synthetic experiments, additional dimensions stand for (artificial) noise variables, put on top of the design and contextual variables.
@@
EggHolder 2 + 4 1 1
Hartmann4D 4 + 3 2 2
Hartmann6D 6 + 6 3 3
Ackley 5 + 8 2 3
Figure 2: Benchmark of the different methods. (a) On real-world datasets, SADCBO (red curve with white markers) performs on par with other baselines and is the top performer for the Robot Pushing task. (b) On synthetic functions, SADCBO outperforms other baselines in three cases out of four. (c) Histograms of phase switching criteriong time for SADCBO computed for the Hartmann6D (c.1) and Hartmann4D problems (c.2). (d) Inclusion probability of each contextual variable for SADCBO computed for the Hartmann6D (d.1) and Hartmann4D problems (d.2). Each panel shows the mean ±$\pm 2$ standard error across $N=100$ trials.
Portfolio Optimization 5D. This dataset was first introduced by Cakmak et al. [2020]. The goal is to optimize three design variables, which stand for the hyperparameters of a trading strategy, to maximize return under random environmental conditions. There are two contextual variables, namely: bid–ask spread and the borrowing cost.
Yacht Hydrodynamics 6D. This dataset comes from the UCI Machine Learning Repository [Gerritsma et al., 2013]. The optimization problem is to maximize the residuary resistance per unit weight of displacement of a yacht by controlling its 5-dimensional hull geometry coefficients. Design variables are the first four dimensions of the hull geometry coefficients. The contextual variables are the last hull geometry dimension and the Froude number.
Molecule structure optimization 21D. This computational chemistry example consists of optimizing the bond angles in an alanine molecule to determine the lowest energy conformer, i.e., the structure the molecule will likely take in nature. These problems are complicated by high dimensionality. We consider the Alanine, a molecule with 21 angular variables: 3 key variables based on prior domain knowledge set as design variables, and 18 other angles treated as contextual variables. Molecular energies are calculated with the AMBER forcefield [Case et al., 2023] at each round of BO.
Robot pushing task 14D. We follow Wang et al. [2017] and consider a control parameter tuning problem for robot pushing. This real-world function returns the distance between a designated goal location and two objects being pushed by two robot hands, whose trajectory is determined by 14 parameters specifying the location, rotation, velocity and moving direction, among others. There are 6 design variables and 8 contextual variables.
Synthetic experiments. We also consider four synthetic test functions, (see Tables 2 and E.2 for details). A min-max transformation is performed on the input data, scaling it to the unit cube: $\mathcal{X}\times\mathcal{Z}=[0,1]^{d+c}$. Similarly, the output is scaled between $[0,1]$ and a noise term $\varepsilon\sim\mathcal{N}(0,\sigma^{2}_{\text{noise}})$ is added with $\sigma^{2}_{\text{noise}}=0.001$. The contextual variable distribution is $p(\mathbf{z})=\mathcal{U}([0,1]^{c})$.
@@ 5.2 Results
In each plot from Figure 2, we report the best value found by each baseline as a function of the number of iterations. In real-world experiments (Figure 2a), SADCBO (in red with white markers) quickly converges to the optimum. SADCBO consistently outperforms the first baselines VBO and CUBO, even though in the Molecular Shape example, SADCBO and CUBO perform on par due to the good choices of the domain experts on the design variables. Except for the Robot Pushing task, the difference between SADCBO and CBO (in blue) is marginal in the real-world experiments. The latter enhances the surrogate model with information from sampled contexts, while our method may even optimize selected contextual variables if needed. Given that these baselines perform similarly, combined with the observation that optimizing only design variables (CUBO, in yellow) produces poor results for the Portfolio and Yacht problems, we can conclude that contextual variables play a significant part in maximizing these two objectives. The cost-aware BO baseline CaBO performs poorly in all tasks. Dropout and MMDBO consistently underperforms, except on the Yacht example for the latter. These baselines perform variable selection in a random manner for Dropout and using Hilbert-Schmidt Independence Criterion for MMDBO [Gretton et al., 2007], two strategies that do not seem to surpass the Feature Collapse method implemented in SADCBO. This observation highlights the need for an informed variable selection strategy. In-depth findings for the Molecule experiment are presented in Section E.1 and provide additional explanations as to why SADCBO clearly outperforms MMDBO and Dropout.
@@
Figure 2b displays the best value found by each baseline for synthetic functions. SADCBO ranks first on 3 out of 4 examples, closely followed by the cost-aware baseline CaBO, which performs much better on synthetic experiments than on the real-world ones. The contextual BO baseline CBO that obtained second to best results in real-world experiments, is now less performant, due to the fact that it does not optimize the context, similarly as CUBO. This seems to be particularly critical for Ackley5D, whereas for Hartmann6D/Hartmann4D, simply enhancing the surrogate with contextual variable observation already leads to a large performance gap between CUBO and CBO. Lastly, VBO does a poor job as it optimizes every variable, thus spending a large fraction of the budget every iteration.
For Hartmann6D and Hartmann4D, Figure 2c reports the time at which SADCBO’s switching criterion (Equation 12) kicks in, in proportion to the total budget, demonstrating that both phases are leveraged in our approach.
Finally, Figure 2d reports the sensitivity indices computed at each iteration for each contextual variable, averaged across whole trajectories of multiple trials. For Hartmann6D, the results match the Sobol sensitivity analysis results (Table S1), even though global sensitivity indices may differ from sensitivity indices with respect to the function optimum. Similar findings apply to Hartmann4D (Table S2). Results for other problems can be found in Figure S2.
Main takeaways. Quantitatively, SADCBO achieves the best overall performances, ranking first in 7 out of 8 problems, although other methods obtained comparable performances on 5 out of 8 problems. The second-best and third-best methods, CBO and MMDBO, both severely underperform in two examples (Ackley and Hartmann6 for CBO, Molecular Shape and EggHolder for MMDBO). While the improvements provided by SADCBO may seem marginal, they are consistent across the benchmark. We hypothesize that this consistent behavior stems from our two-stage approach, which allows SADCBO to be versatile. SADCBO can handle both cases where the impact of the contextual variables on the function is limited (hence it is not worth spending budget to control them) and cases where spending budget leads to informative queries are simultaneously well-handled. For instance, SADCBO effectively reverted to a CBO algorithm in the Molecular Shape problem, due to an optimization phase mostly triggered at the end of the run. Meanwhile, for the Ackley function, the optimization phase was triggered in the first quarter of the budget on average, leading to SADCBO outperforming CBO.